# A Prefrontal Cortex-inspired Architecture for Planning in Large Language Models

### v3

@@ Abstract
Large language models (LLMs) demonstrate impressive performance on a wide variety of tasks, but they often struggle with tasks that require multi-step reasoning or goal-directed planning. To address this, we take inspiration from the human brain, in which planning is accomplished via the recurrent interaction of specialized modules in the prefrontal cortex (PFC). These modules perform functions such as conflict monitoring, state prediction, state evaluation, task decomposition, and task coordination. We find that LLMs are sometimes capable of carrying out these functions in isolation, but struggle to autonomously coordinate them in the service of a goal. Therefore, we propose a black box architecture with multiple LLM-based (GPT-4) modules. The architecture improves planning through the interaction of specialized PFC-inspired modules that break down a larger problem into multiple brief automated calls to the LLM. We evaluate the combined architecture on three challenging planning tasks – graph traversal, Tower of Hanoi, and logistics – finding that it yields significant improvements over both standard LLM methods (zero-shot prompting, in-context learning) and competitive baselines (chain-of-thought, multi-agent debate). These results demonstrate the benefit of utilizing knowledge from cognitive neuroscience to improve planning in LLMs.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) (Devlin et al., 2019; Brown et al., 2020) have recently emerged as highly capable generalist systems with a surprising range of emergent capacities (Srivastava et al., 2022; Wei et al., 2022a; Webb et al., 2023). They have also sparked broad controversy, with some suggesting that they are approaching general intelligence (Bubeck et al., 2023), and others noting a number of significant deficiencies (Mahowald et al., 2023). A particularly notable shortcoming is their poor ability to plan or perform faithful multi-step reasoning (Valmeekam et al., 2023; Dziri et al., 2023). Recent work (Momennejad et al., 2023) has evaluated the extent to which LLMs might possess an emergent capacity for planning and exploiting cognitive maps, the relational structures that humans and other animals utilize to perform planning (Tolman, 1948; Tavares et al., 2015; Behrens et al., 2018). This work found that a variety of LLMs, ranging from small, open-source models (e.g., LLaMA-13B and Alpaca-7B) to large, state-of-the-art models (e.g., GPT-4), displayed systematic shortcomings in planning tasks that suggested an inability to reason about cognitive maps. Common failure modes included a tendency to ‘hallucinate’ (e.g., to imagine non-existent paths), and to fall into loops. This work raises the question of how LLMs might be improved so as to enable a capacity for planning.
In the present work, we take a step toward improving planning in LLMs, by taking inspiration from the planning mechanisms employed by the human brain. Planning is generally thought to depend on the prefrontal cortex (PFC) (Owen, 1997; Momennejad et al., 2018; Momennejad, 2020; Russin et al., 2020; Brunec & Momennejad, 2022; Mattar & Lengyel, 2022), a region in the frontal lobe that is broadly involved in executive function, decision-making, and reasoning (Miller & Cohen, 2001). Research in cognitive neuroscience has revealed the presence of several subregions or modules within the PFC that appear to be specialized to perform certain functions. These include functions such as conflict monitoring (Botvinick et al., 1999); state prediction and state evaluation (Wallis, 2007; Schuck et al., 2016); and task decomposition and task coordination  (Ramnani & Owen, 2004; Momennejad & Haynes, 2012, 2013). Human planning then emerges through the coordinated and recurrent interactions among these specialized PFC modules, rather than through the activity of a single, monolithic system.
Figure 1: LLM-PFC architecture. The agent receives states from the environment and high-level goals. These are processed by a set of specialized LLM modules. The $\operatorname{TaskDecomposer}$ receives high-level goals and generates a series of subgoals. The $\operatorname{Actor}$ generates proposed actions given a state and a subgoal. The $\operatorname{Monitor}$ gates these proposed actions based on whether they violate certain constraints (e.g., task rules) and provides feedback to the $\operatorname{Actor}$. The $\operatorname{Predictor}$ predicts the next state given the current state and a proposed action. The $\operatorname{Evaluator}$ is used to estimate the value of a predicted state. The $\operatorname{Predictor}$ and $\operatorname{Evaluator}$ are used together to perform tree search. The $\operatorname{Orchestrator}$ determines when each subgoal has been achieved, and when the final goal has been achieved, at which point the plan is emitted to the environment as a series of actions. These modules are inspired by the specific PFC subregions depicted on the left.
@@
An interesting observation is that LLMs often seem to display some of these capacities when probed in isolation, even though they are unable to reliably integrate and deploy these capacities in the service of a goal. For instance Momennejad et al. (2023) noted that LLMs often attempt to traverse invalid or hallucinated paths in planning problems (e.g., to move between rooms that are not connected), even though they can correctly identify these paths as invalid when probed separately. This suggests the possibility of a PFC-inspired approach, in which planning is carried out through the coordinated activity of multiple LLM modules, each of which is specialized to perform a distinct process.
With this goal in mind, we propose LLM-PFC (Figure 1), an architecture composed of modules that are specialized to perform specific PFC-inspired functions. Each module consists of an LLM instance (GPT-4), constructed through a combination of prompting and few-shot in-context learning. We specifically propose modules that perform error monitoring, action proposal, state prediction, state evaluation, task decomposition, and task coordination. It is suggested that the coordinated activity of multiple PFC subregions performs tree search during planning (Owen, 1997; Daw et al., 2005; Wunderlich et al., 2012; Doll et al., 2015). Thus, our approach combines action proposal, state prediction, and state evaluation to perform tree search.
@@
We evaluate LLM-PFC on three challenging planning tasks. First, we performed controlled experiments on a set of graph traversal tasks using the CogEval protocol  (Momennejad et al., 2023). These tasks require navigation in novel environments based on natural language descriptions, and have been shown to be extremely challenging for LLMs, including GPT-4. Second, we investigate Tower of Hanoi (ToH), a classic problem solving task that requires multi-step planning (Simon, 1975), and for which performance is known to be heavily dependent on PFC function (Goel & Grafman, 1995; Fincham et al., 2002). Finally, we investigate a more complex, real-world planning task involving logistics (transportation of goods) (Valmeekam et al., 2023). We find that our approach significantly improves LLM performance on all three tasks. Ablation experiments further indicate that each of the individual modules plays an important role in the overall architecture’s performance. Taken together, these results indicate the potential of a PFC-inspired approach to improve the reasoning and planning capabilities of LLMs.
@@ 2 Approach
The LLM-PFC architecture is constructed from a set of a specialized LLM modules, each of which performs a specific PFC-inspired function. In the following sections, we first describe the functions performed by each module, and then describe how they interact to generate a plan.
Function $\operatorname{ProposeAction}$($x,y,B$):
       $\sigma\leftarrow$ false // Initialize validity
$E\leftarrow$
[…]
while $\sigma$ is false do
             $A\leftarrow\operatorname{Actor}(x,y,E,B)$ // Sample B actions
$\sigma,\epsilon\leftarrow\operatorname{Monitor}(x,A)$
// Determine validity and provide feedback
$E\leftarrow E\cup\{\epsilon\}$
// Accumulate feedback
       end while
return $A$
Algorithm 1 Action proposal loop. $\operatorname{ProposeAction}$ takes a state $x$ and a goal $y$ and generates $B$ potential actions $A=a_{b=1}\dots a_{b=B}$. This is implemented via a loop, in which the $\operatorname{Actor}$ first proposes potential actions, and the $\operatorname{Monitor}$ then assesses those actions according to certain constraints (e.g., task rules), providing feedback if any of the actions are deemed to be invalid. This continues until the proposed actions are considered valid. See Sections A.6.2 and A.6.3 for more details.
@@ 2.1 Modules
LLM-PFC contains the following specialized modules, each constructed from a separate LLM instance through a combination of prompting and few-shot ($\leq 3$ examples) in-context learning (described in greater detail in section A.6):
$\bullet$ $\operatorname{\mathbf{TaskDecomposer}}$. The $\operatorname{TaskDecomposer}$ receives the current state $x$ and a goal $y$ and generates a set of subgoals $Z$ that will allow the agent to gradually work toward its final goal. This module is inspired by the anterior PFC (aPFC), which is known to play a key role in task decomposition through the generation and maintenance of subgoals (Ramnani & Owen, 2004). In the present work, the $\operatorname{TaskDecomposer}$ is only utilized to generate a single intermediate goal, though in future work we envision that it will be useful to generate a series of multiple subgoals.
$\bullet$ $\operatorname{\mathbf{Actor}}$. The $\operatorname{Actor}$ receives the current state $x$ and a subgoal $z$ and proposes $B$ potential actions $A=a_{b=1}\dots a_{b=B}$. The $\operatorname{Actor}$ can also receive feedback $\epsilon$ from the $\operatorname{Monitor}$ about its proposed actions. This module can be viewed as being analogous to the dorsolateral PFC (dlPFC) which plays a role in decision making through top-down control and guidance of lower-order premotor and motor regions (Miller & Cohen, 2001).
$\bullet$ $\operatorname{\mathbf{Monitor}}$. The $\operatorname{Monitor}$ assesses the actions proposed by the $\operatorname{Actor}$ to determine whether they are valid (e.g., whether they violate the rules of a task). It emits an assessment of validity $\sigma$, and also feedback $\epsilon$ in the event the action is deemed invalid. This module is inspired by the Anterior Cingulate Cortex (ACC), which is known to play a role in conflict monitoring (Botvinick et al., 1999), i.e., detecting errors or instances of ambiguity.
$\bullet$ $\operatorname{\mathbf{Predictor}}$. The $\operatorname{Predictor}$ receives the current state $x$ and a proposed action $a$ and predicts the resulting next state $\tilde{x}$. The $\operatorname{Predictor}$ is inspired by the Orbitofrontal cortex (OFC), which plays a role in estimating and predicting task states. In particular, it has been proposed that the OFC plays a key role in encoding cognitive maps: representations of task-relevant states and their relationships to one another (Schuck et al., 2016).
$\bullet$ $\operatorname{\mathbf{Evaluator}}$. The $\operatorname{Evaluator}$ receives a next-state prediction $\tilde{x}$ and produces an estimate of its value $v$ in the context of goal $y$. This is accomplished by prompting the $\operatorname{Evaluator}$ (and demonstrating via a few in-context examples) to estimate the minimum number of steps required to reach the goal (or subgoal) from the current state. The $\operatorname{Evaluator}$ is also inspired by the OFC which, in addition to predicting task states, plays a key role in estimating the motivational value of those states (Wallis, 2007).
$\bullet$ $\operatorname{\mathbf{Orchestrator}}$. The $\operatorname{Orchestrator}$ receives the current state $x$ and a subgoal $z$ and emits an assessment $\Omega$ of whether the subgoal has been achieved. When the $\operatorname{Orchestrator}$ determines that all subgoals (including the final goal) have been achieved, the plan is emitted to the environment as a series of actions. This module is also inspired by the aPFC, which is thought to both identify subgoals and coordinate their sequential execution (Ramnani & Owen, 2004).
@@ 2.2 Action Proposal Loop
The $\operatorname{Actor}$ and $\operatorname{Monitor}$ interact via the $\operatorname{ProposeAction}$ function (Algorithm 1). The $\operatorname{Actor}$ proposes actions which are then gated by the $\operatorname{Monitor}$. If the $\operatorname{Monitor}$ determines that the actions are invalid (e.g., they violate the rules of a task), feedback is provided to the $\operatorname{Actor}$, which then proposes an alternative action. In the brain, a similar process is carried out by interactions between the ACC and dorsolateral PFC (dlPFC). The ACC is thought to recruit the dlPFC under conditions of conflict (e.g., errors or ambiguity), which then acts to resolve the conflict through top-down projections to lower-order control structures (e.g., premotor and motor cortices) (Miller & Cohen, 2001; Shenhav et al., 2013).
Function $\operatorname{Search}$($l,L,B,x,y$):
       $V_{l}\leftarrow\{\}$ // Initialize value record
$\tilde{X}_{l}\leftarrow\{\}$
// Initialize next-state record
$A_{l}\leftarrow\operatorname{ProposeAction}(x,y,B)$
// Propose B actions
for $b$ in $1\ldots B$ do
$\tilde{X}_{l}\leftarrow\tilde{X}_{l}\cup\{\tilde{x}_{lb}\}$
// Update next-state record
$\Omega\leftarrow\operatorname{Orchestrator}(\tilde{x}_{lb},y)$
// Terminate search if goal achieved
if $l<L$ and $\Omega$ is false then
$V_{l}\leftarrow V_{l}\cup\{v_{l+1}\}$
// Update value record
            else
$V_{l}\leftarrow V_{l}\cup\{v_{lb}\}$
// Update value record
             end if
       end for
 $a_{l}\leftarrow A_{l\operatorname{argmax}(V_{l})}$ // Select action
$\tilde{x}_{l}\leftarrow\tilde{X}_{l{\operatorname{argmax}(V_{l})}}$
// Predicted next-state
return $a_{l},\tilde{x}_{l},v_{l}$
2.3 Search Loop
$\operatorname{ProposeAction}$ is further embedded in a $\operatorname{Search}$ loop (Algorithm 2). The actions emitted by $\operatorname{ProposeAction}$ are passed to the $\operatorname{Predictor}$, which predicts the states that will result from these actions. A limited tree search is then performed, starting from the current state, and then exploring $B$ branches recursively to a depth of $L$ layers. Values are assigned to the terminal states of this search by the $\operatorname{Evaluator}$, and the action leading to the most valuable predicted state is selected. This approach mirrors that of the human brain, in which search is thought to be carried out through the coordinated activity of multiple regions within the PFC, including dlPFC, ACC, and OFC (Owen, 1997; Mattar & Lengyel, 2022).
Function $\operatorname{LLM-PFC}$($x,y,T,L,B$):
@@ 2.3 Search Loop
$Z\leftarrow\operatorname{TaskDecomposer}(x,y)$
// Generate subgoals
@@
$P\leftarrow P$
.append$(a)$ // Update plan
$\Omega\leftarrow\operatorname{Orchestrator}(x,z)$
// Determine if subgoal is achieved
@@
Algorithm 3 LLM-PFC. $\operatorname{LLM-PFC}$ takes a state $x$ and a goal $y$ and generates a plan $P$, a series of actions with a maximum length of $T$. The $\operatorname{TaskDecomposer}$ first generates a set of subgoals $Z$. The agent then pursues each individual subgoal $z$ in sequence, followed by the final goal $y$. At each time step, $\operatorname{Search}$ is called to generate an action and a predicted next-state. Actions are added to the plan until the $\operatorname{Orchestrator}$ determines that the goal has been achieved, or the plan reaches the maximum length $T$.

### v4

# Improving Planning with Large Language Models: A Modular Agentic Architecture

@@ Abstract
Large language models (LLMs) demonstrate impressive performance on a wide variety of tasks, but they often struggle with tasks that require multi-step reasoning or goal-directed planning. Both cognitive neuroscience and reinforcement learning (RL) have proposed a number of interacting functional components that together implement search and evaluation in multi-step decision making. These components include conflict monitoring, state prediction, state evaluation, task decomposition, and orchestration. To improve planning with LLMs, we propose an agentic architecture, the Modular Agentic Planner (MAP), in which planning is accomplished via the recurrent interaction of the specialized modules mentioned above, each implemented using an LLM. MAP improves planning through the interaction of specialized modules that break down a larger problem into multiple brief automated calls to the LLM. We evaluate MAP on three challenging planning tasks – graph traversal, Tower of Hanoi, and the PlanBench benchmark – as well as an NLP task requiring multi-step reasoning (strategyQA). We find that MAP yields significant improvements over both standard LLM methods (zero-shot prompting, in-context learning) and competitive baselines (chain-of-thought, multi-agent debate, and tree-of-thought), can be effectively combined with smaller and more cost-efficient LLMs (Llama3-70B), and displays superior transfer across tasks. These results suggest the benefit of a modular and multi-agent approach to planning with LLMs.
**footnotetext: Equal contribution. Co-first author order is arbitrary and may be swapped when citing this work.$\dagger$\$$\dagger$\$footnotetext: Work performed during internship at Microsoft Research.
@@ 1 Introduction
Large Language Models (LLMs) (Devlin et al., 2019; Brown et al., 2020) have become widely accepted as highly capable generalist systems with a surprising range of emergent capacities (Srivastava et al., 2022; Wei et al., 2022a; Webb et al., 2023). They have also sparked broad controversy, with some suggesting that they are approaching general intelligence (Bubeck et al., 2023), and others noting a number of significant deficiencies (Mahowald et al., 2023). A particularly notable shortcoming is their poor ability to plan or perform faithful multi-step reasoning (Valmeekam et al., 2023; Dziri et al., 2023). Recent work (Momennejad et al., 2023) has evaluated the extent to which LLMs might possess an emergent capacity for planning and exploiting cognitive maps, the relational structures that humans and other animals utilize to perform planning (Tolman, 1948; Tavares et al., 2015; Behrens et al., 2018). This work found that LLMs displayed systematic shortcomings in planning tasks that suggested an inability to reason about cognitive maps. Common failure modes included a tendency to ‘hallucinate’ (e.g., to use non-existent transitions and paths), and to fall into loops. This work raises the question of how LLMs might be improved so as to enable a capacity for planning, especially given the ubiquity of sequential decision making, reasoning, and planning problems across the wide application of generative AI and LLMs.
@@
Here, we take a step toward improving planning with LLMs, by taking inspiration from both cognitive neuroscience and formal theories of decision-making and planning. In traditional theories of planning, such as those found in the field of reinforcement learning (RL) (Sutton & Barto, 2018), planning is carried out via the interaction of several specialized functions or modules, rather than through the activity of a single, monolithic system. For instance, many approaches involve distinct functions for action proposal, state evaluation, subgoal identification, or state prediction, many of which have also been related to the function of specific brain regions (see Section 6 for discussion). An interesting observation is that LLMs are often able to carry out these functions when probed in isolation, but are unable to reliably integrate and orchestrate these capacities in the service of a goal. For instance, Momennejad et al. (2023) noted that LLMs often attempt to traverse invalid or hallucinated paths in planning problems (e.g., to move between rooms that are not connected), even though they can correctly identify these paths as invalid when probed separately. This suggests the possibility of an agentic approach using LLMs, in which planning is carried out through the coordinated and recurrent interaction of multiple LLM modules, each of which is specialized to perform a distinct process.
Figure 1: Modular Agentic Planner (MAP). The agent receives states from the environment and high-level goals. These are processed by a set of specialized LLM modules. The $\operatorname{TaskDecomposer}$ receives high-level goals and generates a series of subgoals. The $\operatorname{Actor}$ generates proposed actions given a state and a subgoal. The $\operatorname{Monitor}$ gates these proposed actions based on whether they violate certain constraints (e.g., task rules) and provides feedback to the $\operatorname{Actor}$. The $\operatorname{Predictor}$ predicts the next state given the current state and a proposed action. The $\operatorname{Evaluator}$ is used to estimate the value of a predicted state. The $\operatorname{Predictor}$ and $\operatorname{Evaluator}$ are used together to perform tree search. The $\operatorname{Orchestrator}$ determines when each subgoal has been achieved, and when the final goal has been achieved, at which point the plan is emitted to the environment as a series of actions.
With this goal in mind, we propose the Modular Agentic Planner (MAP) (Figure 1), an agentic architecture composed of modules that are specialized to perform specific functions within the planning process. Specifically, we have identified and implemented the following key modules: error monitoring, action proposal, state prediction, state evaluation, task decomposition, and task coordination. Action proposal, state prediction, and state evaluation are further combined to perform tree search. All modules are implemented using an LLM, which receives instructions describing the module’s role via prompting and few-shot in-context learning (ICL). The resulting MAP algorithm (Algorithm 1) is implemented via the recurrent interaction of these modules, combining the strengths of classical planning and search algorithms with the use of LLMs as general-purpose world models and planning functions.
@@
We evaluate MAP on four challenging decision-making tasks that require planning and multi-step reasoning. First, we performed controlled experiments on a set of graph traversal tasks according to the CogEval protocol  (Momennejad et al., 2023). These tasks require goal-directed navigation in novel environments (MDPs) described in natural language, of which we selected an environment that was most challenging for LLMs, including GPT-4. Second, we investigate Tower of Hanoi (ToH), a classic problem solving task that requires multi-step planning (Simon, 1975). Third, we investigate the two most challenging tasks in the PlanBench benchmark: mystery BlocksWorld and Logistics (Valmeekam et al., 2023). Finally, we investigate a challenging NLP task that requires multi-step reasoning, StrategyQA (Geva et al., 2021). We find that, when implemented with GPT-4, MAP significantly improves performance on all four tasks (Figures 2 and 3, Tables $\dagger$2 and $\dagger$2), and that the approach can also be effectively implemented with a smaller and more cost-efficient LLM (Llama3-70B, Table 10). Transfer experiments further indicate that MAP displays an improved ability to generalize between tasks, and ablation experiments indicate that each of the individual modules plays an important role in the overall architecture’s performance (Figure 3). Taken together, these results indicate the potential of a modular agentic approach to improve the reasoning and planning capabilities of LLMs.
@@ 2 Approach
What follows first describes the functions performed by each MAP module, and then how they interact to generate a plan.
@@ 2.1 Modules
MAP contains the following specialized modules, each constructed from a separate LLM instance through a combination of prompting and few-shot ($\leq 3$ examples) in-context learning (described in greater detail in section A.7):
$\bullet$ $\operatorname{\mathbf{TaskDecomposer}}$. The $\operatorname{TaskDecomposer}$ receives the current state $x$ and a goal $y$ and generates a set of subgoals $Z$ that will allow the agent to gradually work toward its final goal. In the present work, the $\operatorname{TaskDecomposer}$ is only utilized to generate a single intermediate goal, though in future work we envision that it will be useful to generate a series of multiple subgoals.
$\bullet$ $\operatorname{\mathbf{Actor}}$. The $\operatorname{Actor}$ receives the current state $x$ and a subgoal $z$ and proposes $B$ potential actions $A=a_{b=1}\dots a_{b=B}$. The $\operatorname{Actor}$ can also receive feedback $\epsilon$ from the $\operatorname{Monitor}$ about its proposed actions.
$\bullet$ $\operatorname{\mathbf{Monitor}}$. The $\operatorname{Monitor}$ gates the actions proposed by the $\operatorname{Actor}$ based on their validity (e.g., whether they violate the rules of a task). It emits an assessment of validity $\sigma$, and also feedback $\epsilon$ in the event the action is deemed invalid.
$\bullet$ $\operatorname{\mathbf{Predictor}}$. The $\operatorname{Predictor}$ receives the current state $x$ and a proposed action $a$ and predicts the resulting next state $\tilde{x}$.
$\bullet$ $\operatorname{\mathbf{Evaluator}}$. The $\operatorname{Evaluator}$ receives a next-state prediction $\tilde{x}$ and produces an estimate of its value $v$ in the context of goal $y$. This is accomplished by prompting the $\operatorname{Evaluator}$ (and demonstrating via a few in-context examples) to estimate the minimum number of steps required to reach the goal (or subgoal) from the current state.
$\bullet$ $\operatorname{\mathbf{Orchestrator}}$. The $\operatorname{Orchestrator}$ receives the current state $x$ and a subgoal $z$ and emits an assessment $\Omega$ of whether the subgoal has been achieved. When the $\operatorname{Orchestrator}$ determines that all subgoals (including the final goal) have been achieved, the plan is emitted to the environment as a series of actions.
@@ 2.2 Action Proposal Loop
This section describes MAP’s algorithms, the first of which is the action proposal loop. The $\operatorname{Actor}$ and $\operatorname{Monitor}$ interact via the $\operatorname{ProposeAction}$ function (Supplementary Algorithm 2). The $\operatorname{Actor}$ proposes actions which are then gated by the $\operatorname{Monitor}$. If the $\operatorname{Monitor}$ determines that the actions are invalid (e.g., they violate the rules of a task), feedback is provided to the $\operatorname{Actor}$, which then proposes an alternative action.
2.3 Tree Search
$\operatorname{ProposeAction}$ is further embedded in a $\operatorname{Search}$ loop (Supplementary Algorithm 3). The actions emitted by $\operatorname{ProposeAction}$ are passed to the $\operatorname{Predictor}$, which predicts the states that will result from these actions. A limited tree search is then performed, starting from the current state, and then exploring $B$ branches recursively to a depth of $L$ layers. Values are assigned to the terminal states of this search by the $\operatorname{Evaluator}$, and the action leading to the most valuable predicted state is selected.
Function $\operatorname{MAP}$($x,y,T,L,B$):
@@ 2.3 Tree Search
       $Z\leftarrow\operatorname{TaskDecomposer}(x,y)$ // Generate subgoals
@@
                   $P\leftarrow P$.append$(a)$ // Update plan
                   $\Omega\leftarrow\operatorname{Orchestrator}(x,z)$ // Determine if subgoal is achieved
@@
Algorithm 1 Modular Agentic Planner (MAP). $\operatorname{MAP}$ takes a state $x$ and a goal $y$ and generates a plan $P$, a series of actions with a maximum length of $T$. The $\operatorname{TaskDecomposer}$ first generates a set of subgoals $Z$. The agent then pursues each individual subgoal $z$ in sequence, followed by the final goal $y$. At each time step, $\operatorname{Search}$ (Algorithm 3) is called to generate an action and a predicted next-state. Actions are added to the plan until the $\operatorname{Orchestrator}$ determines that the goal has been achieved, or the plan reaches the maximum length $T$.